Oreo Inc. is a cookie manufacturer making delicious chocolate cookies. Early in 2019, in an effort to produce healthier produWhat is the 2019 Cost of Goods Manufactured for Spinach Oreos

What is the 2019 Cost of Goods Sold for Spinach Oreos

What is the Income from Operations for the Spinach Oreo division

Based on the Income from Operations, was Spinach Oreos a good idea (yes or no)
Answer #1
1) Cost of Production in 2019
Direct material cost per unit
Chocolate                     5.00
Sugar                     3.00
Spinach                     0.30
                    8.30
Direct Labour Cost per unit                     0.45
Factory Overhead = 1,250 per machine hour
Production per mahine hour = 200 cookies
Machine hour to produce 750,000 units = 750,000/200
= 3750 machhine hour
Factory overhead for 750,000 unit               = 3750 X 1,250
4,687,500
Cost of Production in 2019                       = Direct material + Direct Labour + Factory overhaed
Direct Material cost for 750,000 units       = 750,000 X 8.3
62,25,000
Direct Labour Cost = 750,000 X .45
= 3,37,500
Factory overhead = 46,87,500
Total Cost of Production 6,225,000 + 3,37500 + 4,687,500
Cost of production for 750,000 units = 1,12,50,000
Per unit Cost of Spinach Oreos = Total Cost of production/Number of units manufactured
= 11,250,2000 /750,000
= 15 per unit
2) Cost of Godds sold in 2019
Number of Units Sold 730,000 units
Cost goods Sold Cost of goos manufactured + selling and distribution expense
Cost of goods manuactured for 730,000 = 730,000 X 15
10,950,000
Commission 30 % of whole sale price
Whole sale price 25
Commission for 730,000 units 730,000 X (25X30%)
= 730,000 X 7.5
= 5,475,000
Promotional expense = 900,000
Cost of Goods sold = 10,950,000 + 5,475,000 + 900,000 + 1,000,000
= 18,325,000
3) Income from operation of Spinach Oreos
Income from operation = Sales amount - Cost of Goods sold
Sales amount for 730,000 units = 730,000 X 25
= 18,250,000
Cost of Goods sold = 18,325,000
Income from operation (Loss) = 18,250,000 - 18,325,000
= (75,000)
4) Based on income from operatoin Spinach Oreos is good idea
Yes
In the year 2019 the loss is only $ 75,000 because of the aditional promotional expense incurred
Next year ownwards they are not required any additional promotional expense in the market